At Jerry's birth, his mother was 27 years old. At what age will Jerry be 40 percent of his mother's age
Answer by macston(5194)

.
x=age at which Jerry will be 40% mothers age
.
0.4(27+x)=x
10.8+0.4x=x
10.8=0.6x
18=x
ANSWER: When Jerry is 18 he will be 40% of his mother's age.
.
CHECK:
When Jerry is 18, his mother will be 27+18=45.
.
0.40(45)=18
18=18